# Multi-Partite Output Regulation of Multi-Agent Systems

### v1

@@ Abstract
This article proposes a simple, graph-independent perspective on partitioning the node set of a graph and provides multi-agent systems (MASs) with objectives beyond cooperation and bipartition. Specifically, we first introduce the notion of $k$-partition transformation to achieve any desired partition of the nodes. Then, we use this notion to formulate the multi-partite output regulation problem (MORP) of heterogeneous linear MASs, which comprises the existing cooperative output regulation problem (CORP) and bipartite output regulation problem (BORP) as subcases. The goal of the MORP is to design a distributed control law such that each follower that belongs to the same set in the partition asymptotically tracks a predefined multiple of a reference while ensuring the internal stability of the closed-loop system. It is shown that the necessary and sufficient conditions for the solvability of the MORP with a feedforward-based distributed control law follow from the CORP and lead to the first design strategy for the control parameters. However, it has a drawback in terms of scalability due to a partition-dependent condition. We prove that this condition is implied by its partition-independent version under a mild structural condition. This implication yields the second design strategy that is much more scalable than the first one. Finally, numerical examples are provided to illustrate the generality of the MORP and compare both design strategies regarding scalability.
@@ I-A Motivation and Literature Review
Considering the studies on distributed control of multi-agent systems (MASs), two main frameworks become distinguishable: cooperative [1] and bipartite [2, 3]. In both frameworks, consensus, synchronization, leader-following consensus, formation, and output regulation stand out as the most recognized problems. Regardless of the problem, MASs have a common objective in the cooperative framework, whereas they potentially have two opposed objectives in the bipartite framework.
@@
This article is mainly motivated by the need for a flexible framework that extends beyond cooperation and bipartition within MASs to accommodate multiple, shifting mission objectives, particularly in adverse operating environments like suppression and destruction of enemy air defense operations [4]. Given that the output regulation problems for MASs consider heterogeneity in agent dynamics, pave the way for a capability of tracking and rejecting a large class of signals, and contain problems such as leader-following consensus and formation, we propose the multi-partite output regulation problem (MORP) to enhance tactical flexibility. The MORP includes the cooperative output regulation problem (CORP) and bipartite output regulation problem (BORP) as subcases.
@@
Analogous to the output regulation problem, the CORP has been mainly treated using feedforward [5, 6, 7, 8, 9] and internal model [10, 11, 12, 13, 14, 15] approaches. In the former, the feedforward gain of each agent is based on the solution of the regulator equations, which are linear matrix equations (LMEs) determined by the exosystem and agent dynamics, making it not robust to parameter uncertainties. While the latter is known to be robust against small parameter variations, it cannot be applied when the transmission zeros condition does not hold.
@@
The studies with the feedforward approach can be divided into two: (i) [5, 6, 7, 8] use virtual information exchange (i.e., controller states of neighboring agents); (ii) [9] uses only physical information exchange (i.e., relative outputs of neighboring agents). Similarly, one can classify the articles solving the CORP via the internal model approach into two: (i) [16, 10, 11] use virtual information exchange; (ii) [12, 13, 14, 15] use only physical information exchange. Compared to the CORP, fewer studies exist on the BORP. It is tackled using the feedforward approach with virtual information exchange in [17, 18, 19] and the internal model approach with physical information exchange in [20, 3]. In the CORP (BORP), the feedforward approach with virtual information exchange uses a distributed observer to provide the estimated state (estimated state or its additive inverse) of the exosystem to every agent.
Apart from the bipartite framework, there have been efforts to extend the number of feasible objectives of MASs. The notable ones are cluster consensus [21, 22, 23], scaled consensus [24], and kernel manipulation of the Laplacian matrix [25, 26]. However, those studies are limited to consensus problems over single-integrator agent dynamics.
@@ I-B Contribution
This article formulates a general distributed control problem for heterogeneous MASs. To solve this problem, it provides two design strategies for a feedforward-based distributed control law relying on virtual information exchange with comparable advantages and disadvantages.
While the partition in the bipartite framework is graph-dependent, we first introduce the notion of $k$-partition transformation to render the partition of the node set independent of the given graph. Hence, the node set can now admit up to $N$-partition, where $N$ is the cardinality of the node set. In fact, there are infinitely many $k$-partition transformations so that the node set can be partitioned the Bell number of ways.
Then, leveraging this notion, we formulate the MORP, which includes the CORP and BORP as special cases (see Remark 2 and Examples 1 and 2). To this end, we consider a heterogeneous MAS in the form
@@
This autonomous linear system yields the reference (i.e., $-F_{i}v$) and disturbance (i.e., $E_{i}v$) signals for the MAS (I-B). The goal of the MORP is to design a distributed control law such that each subsystem in the same set within the partition asymptotically tracks a common predefined multiple of the reference (i.e., a predefined $k$-partition transformation term multiple of the reference) and rejects the disturbance while ensuring the internal stability of the closed-loop system.
The key observation leading to necessary and sufficient conditions for the solvability of the MORP with a feedforward-based distributed control law is that the MORP can be realized as the CORP after redefining the reference. This results in the first design strategy for each subsystem’s control parameters. Though straightforward, this strategy requires each subsystem to recompute a solution for the regulator equations each time the $k$-partition transformation term is updated. An intriguing question arises from this drawback: Is it possible to have a design strategy that eliminates the recomputation of a solution for the regulator equations whenever the $k$-partition transformation term changes? The proposed second design strategy gives an affirmative answer under a mild structural condition, which is independent of the $k$-partition transformation term. Therefore, this design strategy is highly scalable for a set of $k$-partition transformations compared to the first one.
We compare both design strategies in two categories.
  1. 1.
    Scalability: Given a finite set of $k$-partition transformation terms with $M$ elements, each subsystem employing the first design strategy computes $M$ solution pairs to the corresponding regulator equations. On the other hand, the second one makes the regulator equations independent of the $k$-partition transformation term at the cost of adding an LME dependent only on the subsystem parameters. Each subsystem computes a solution pair to these regulator equations and a solution to the LME only once. In other words, regardless of the size of the set of $k$-partition transformation terms, even infinity, they are solved once. Thus, this strategy is much more scalable than the first one (see Remark 12 and Example 3).
  2. 2.
    Conservatism: The second design strategy is applicable only when subsystems are subject to matched disturbances. Though this structural condition is quite common (e.g., see unmanned aerial vehicles (UAVs) [27] and nonholonomic wheeled robots [28]), the first design strategy can be applied even if this condition does not hold (see Remark 13).
@@ I-C Notation
The real part of a complex number $\lambda$ is denoted by $\mathrm{Re}(\lambda)$. The closed right (left) half complex plane is denoted by $\mathrm{CRHP}$ ($\mathrm{CLHP}$). The open right half complex plane is denoted by $\mathrm{ORHP}$. We write $I_{n}$ for the $n\times n$ identity matrix and $\mathrm{diag}(w_{1},\ldots,w_{n})$ for a diagonal matrix with scalar entries $w_{1},\ldots,w_{n}$ on its diagonal. The spectrum of a square matrix $X\in\mathbb{R}^{n\times n}$ is denoted by $\mathrm{spec}(X)$. The matrix obtained by replacing each entry of $X$ with its absolute value is denoted by $|X|$. The image of a matrix $Z\in\mathbb{R}^{n\times m}$ is denoted by $\mathrm{im}\hskip 0.85355ptZ$. The vector formed by stacking columns of $Z$ is denoted by $\mathrm{vec}\hskip 0.85355ptZ$. The Kronecker product is denoted by $\otimes$.
@@ II-A Arbitrary Partition of Nodes
A matrix $S=\mathrm{diag}(s_{1},\ldots,s_{N})\in\mathbb{R}^{N\times N}$ with exactly $1\leq k\leq N$ distinct entries is called a $k$-partition transformation. A 1-partition or 2-partition transformation $S$ is a gauge transformation if $s_{i}\in\{-1,1\}$ for $i=1,\ldots,N$.
Given a signed directed graph, the bipartite framework partitions the nodes according to the signs of the adjacency matrix entries (i.e., the signs of the edge weights). This yields at most $2$-partition of nodes. In contrast, we partition the nodes according to a $k$-partition transformation. Since a $k$-partition transformation can be chosen for any $k\in\{1,\ldots,N\}$, the node set can now admit not only 1-partition or 2-partition but also more than 2-partition. In what follows, we formally discuss the $k$-partition of nodes.
Let $\mathcal{G}(\mathcal{A})$ be a signed directed graph, and let $S$ be a $k$-partition transformation. Then, there exist $k$ positive integers $i_{1},\ldots,i_{k}$ such that $s_{i_{1}},\ldots,s_{i_{k}}$ are $k$ distinct real numbers. For each $p\in\left\{1,\ldots,k\right\}$, let $\mathcal{N}_{p}$ denote the following set
@@
(i) Clearly, $i_{p}\in\mathcal{N}_{p}$. (ii) Let $\mathcal{N}_{p}\in\mathcal{C}$ and $\mathcal{N}_{r}\in\mathcal{C}$ with $p\neq r$, but assume for contradiction that $\mathcal{N}_{p}\cap\mathcal{N}_{r}\neq\emptyset$. Then, let $j\in\mathcal{N}_{p}\cap\mathcal{N}_{r}$; hence, $j\in\mathcal{N}_{p}$ and $j\in\mathcal{N}_{r}$. By definition, $s_{j}=s_{i_{p}}$ and $s_{j}=s_{i_{r}}$. Thus, $s_{i_{p}}=s_{i_{r}}$, which contradicts $s_{i_{p}}$ and $s_{i_{r}}$ being distinct. (iii) Since every $\mathcal{N}_{p}\in\mathcal{C}$ is a subset of $\mathcal{N}$, the inclusion $\bigcup_{\mathcal{N}_{p}\in\mathcal{C}}\mathcal{N}_{p}\subseteq\mathcal{N}$ holds. Let $j\in\mathcal{N}$. Then, there exists a $p\in\left\{1,\ldots,k\right\}$ such that $s_{j}=s_{i_{p}}$ due to the fact that $S$ has exactly $k$ distinct entries. Hence, $j\in\mathcal{N}_{p}$ for some $\mathcal{N}_{p}\in\mathcal{C}$. This proves that the inclusion $\mathcal{N}\subseteq\bigcup_{\mathcal{N}_{p}\in\mathcal{C}}\mathcal{N}_{p}$ holds. ∎
@@
Owing to Definition 2 and Lemma 1, $k$-partition transformations can achieve arbitrary partition of nodes. For example, let $\mathcal{G}(\mathcal{A})$ be a signed directed graph with $5$ nodes. For $k=1,2,3,4,5$, there exist infinitely many $k$-partition transformations that can be used to define $1,15,25,10,1$ number of $k$-partitions of the node set, respectively. These numbers correspond to the Stirling numbers of the second kind for respective values of $k$ (e.g., see Section 5.1 in [29]). Their sum is $52$, which is the associated Bell number.
@@ II-B MORP
As in the context of CORP and BORP, the subsystems of (I-B), considered followers, and the exosystem (2), considered the leader, constitute a MAS of $N+1$ agents. To model the information exchange between $N$ followers, we use a time-invariant signed directed graph ${\mathcal{G}}(\mathcal{A})$ without self-loops as $\mathcal{N}=\left\{1,\ldots,N\right\}$, where node $i\in\mathcal{N}$ corresponds to follower $i$, and for each $j,i\in\mathcal{N}$, we put $(j,i)\in\mathcal{E}$ if, and only if, follower $i$ has access to the information of follower $j$. The leader is included in the information exchange model by augmenting the graph ${\mathcal{G}}(\mathcal{A})$. Specifically, let $\mathcal{G}(\bar{\mathcal{A}})$ be an augmented signed direct graph with $\bar{\mathcal{N}}=\mathcal{N}\cup\{0\}$, $\bar{\mathcal{E}}=\mathcal{E}\cup\mathcal{E}^{\prime}$, where $\mathcal{E}^{\prime}\subseteq\{(0,i)\ |\ i\in\mathcal{N}\}$, and $\bar{\mathcal{A}}\in\mathbb{R}^{(N+1)\times(N+1)}$. Here, node 00 corresponds to the leader and for any $i\in\mathcal{N}$, we put $(0,i)\in\mathcal{E}^{\prime}$ if, and only if, follower $i$ has access to the information of the leader. For any $i\in\mathcal{N}$, the pinning gain $f_{i}>0$ if $(0,i)\in\mathcal{E}^{\prime}$ and $f_{i}=0$ otherwise. The pinning gain matrix is defined by $\mathcal{F}=\mathrm{diag}(f_{1},\ldots,f_{N})$.

### v2

@@ Abstract
This article proposes a simple, graph-independent perspective on partitioning the node set of a graph and provides multi-agent systems (MASs) with objectives beyond cooperation and bipartition. Specifically, we first introduce the notion of $k$-partition transformation to achieve any desired partition of the nodes. Then, we use this notion to formulate the multi-partite output regulation problem (MORP) of heterogeneous linear MASs, which comprises the existing cooperative output regulation problem (CORP) and bipartite output regulation problem (BORP) as subcases. The goal of the MORP is to design a distributed control law such that each follower that belongs to the same set in the partition asymptotically tracks a scalar multiple of the reference while ensuring the internal stability of the closed-loop system. It is shown that the necessary and sufficient conditions for the solvability of the MORP with a feedforward-based distributed control law follow from the CORP and lead to the first design strategy for the control parameters. However, it has a drawback in terms of scalability due to a partition-dependent condition. We prove that this condition is implied by its partition-independent version under a mild structural condition. This implication yields the second design strategy that is much more scalable than the first one. Finally, an experiment is conducted to demonstrate the MORP’s flexibility, and two numerical examples are provided to illustrate its generality and compare both design strategies regarding scalability.
@@ I-A Motivation and Literature Review
Considering the studies on distributed control of multi-agent systems (MASs), two main frameworks become distinguishable: cooperative [1] and bipartite [2],[3]. The most recognized problems of both frameworks include consensus, formation tracking, and output regulation. Regardless of the problem, MASs have a common objective in the cooperative framework, whereas they potentially have two opposed objectives in the bipartite framework.
@@
This article is mainly motivated by the need for a flexible framework that extends beyond cooperation and bipartition within MASs to accommodate multiple, shifting mission objectives in adverse operating environments. For example, consider networked uninhabited aerial vehicles (UAVs) tasked with suppressing and destroying enemy air defense [4]. During this operation, having the flexibility to achieve arbitrarily changing formations is paramount to avoid radars while maximizing damage. Given that the output regulation problems for MASs allow high-order nonidentical agent dynamics and contain typical cooperative control problems such as leader-following consensus and formation tracking, we propose the multi-partite output regulation problem (MORP) to enhance tactical flexibility of MASs. The MORP includes the cooperative output regulation problem (CORP) and the bipartite output regulation problem (BORP) as special cases.
@@
Analogous to the output regulation problem, the CORP has been mainly treated using feedforward[5, 6, 7, 8, 9] and internal model [10, 11, 12, 13, 14] approaches. In the former, the feedforward gain of each agent is based on the solution of the regulator equations, which are linear matrix equations (LMEs) determined by the exosystem and agent dynamics, making it not robust to parameter uncertainties. While the latter is known to be robust against small parameter variations, it cannot be applied when the transmission zeros condition does not hold.
@@
Compared to the CORP, fewer studies exist on the BORP. It is tackled using the feedforward approach in [15, 16, 17] and the internal model approach in [18]. In the CORP (BORP), the feedforward approach, with controller state exchange of neighboring agents, uses a distributed observer to provide the estimated state (estimated state or its additive inverse) of the exosystem to every agent.
Apart from the bipartite framework, there have been efforts to increase the number of objectives in MASs. The notable ones are cluster consensus [19, 20, 21], scaled consensus [22],[23] and kernel manipulation of the Laplacian matrix [24],[25]. Yet, those studies are limited to consensus problems over first or second order agent dynamics. The cluster problem is also extended to high-order heterogeneous MASs [26, 27]. The proposed MORP differs from the cluster consensus in two main aspects: First, the partitioning in the MORP is independent of the underlying graph, allowing each agent to determine its set in the partition, and hence, its objective, whereas the clustering in cluster consensus is graph-dependent, preventing each agent from self-determining its cluster. Second, while the MORP requires only one leader to generate multiple objectives, the cluster consensus requires at least two leaders to yield more than one objective.
@@ I-B Contribution
This article formulates a general distributed control problem for heterogeneous MASs. To solve this problem, it provides two design strategies with comparable advantages and disadvantages for a feedforward-based distributed control law involving a distributed observer.
The partition of the node set of a graph in the bipartite framework is graph-dependent (e.g., see Remark 2.1 and Equation (8) in [18]). In particular, each agent needs to know the graph’s adjacency matrix to determine the set in which it lies in the bipartition, and hence, track the reference or its additive inverse. To render the partition of the node set independent of the graph, we introduce the notion of $k$-partition transformation. This not only allows the node set to admit up to $N$-partition, where $N$ is the cardinality of the node set, but also provides each agent with the flexibility to self-determine the set in which it lies in the $k$-partition, and hence, which scalar multiple of the reference to track.
By leveraging this notion, we formulate the MORP, which includes the CORP and BORP as special cases. To this end, we consider a heterogeneous MAS in the form
@@
This autonomous linear system yields the disturbances $E_{i}v$ and $G_{i}v$ and the reference denoted by $-F_{i}v$ for subsystem $i$. The goal of the MORP is to design a distributed control law such that each subsystem in the same set within the partition asymptotically tracks a common scalar multiple of the reference (i.e., a $k$-partition transformation term multiple of the reference) and rejects the disturbance while ensuring the internal stability of the closed-loop system.
Although the MORP can be cast as the CORP, and hence, the necessary and sufficient solvability conditions of the MORP follow from the CORP, the regulator equations become partition-dependent (see Condition 3). Consequently, the immediate design strategy for control parameters, which is called the first design strategy, requires each subsystem to recompute a solution pair for the regulator equations each time the $k$-partition transformation term is updated. This is a drawback that does not exist in the CORP.
An intriguing question arises from the drawback above: Is it possible to have a design strategy that eliminates the recomputation of a solution for the regulator equations whenever the $k$-partition transformation term changes? Theorem 2 paves the way for an affirmative answer under a mild structural condition. Specifically, it shows that the solvability of the partition-independent regulator equations (3) and the introduced LME (10), which solely depends on the subsystem data, ensures the existence of a solution pair to the partition-dependent regulator equations (3). More importantly, Theorem 2 provides an analytical formula (2) for such a pair in terms of $s_{i}$ and the solutions of (3) and (10). This leads to the second design strategy that does not require recomputing a solution to any equations when the $k$-partition transformation term changes. Accordingly, it is significantly more scalable than the first one. The discussion on the proposed design strategies’ scalability and conservatism is summarized in Table I by referring to the corresponding conditions and results of the paper. An experiment is conducted to demonstrate the MORP’s flexibility in accommodating shifting mission objectives. Two numerical examples are also provided to showcase its generality and compare both design strategies regarding scalability.
@@ I-C Notation
The real part of a complex number $\lambda$ is denoted by $\mathrm{Re}(\lambda)$. The closed right (left) half complex plane is denoted by $\mathrm{CRHP}$ ($\mathrm{CLHP}$). The open right half complex plane is denoted by $\mathrm{ORHP}$. We write $I_{n}$ for the $n\times n$ identity matrix, $0_{n\times m}$ or 00 for the $n\times m$ zero matrix, $\mathrm{diag}(w_{1},\ldots,w_{n})$ for the diagonal matrix with scalar entries $w_{1},\ldots,w_{n}$ on its diagonal, and $\otimes$ for the Kronecker product. The spectrum of a square matrix $X\in\mathbb{R}^{n\times n}$ is denoted by $\mathrm{spec}(X)$. The matrix obtained by replacing each entry of $X$ with its absolute value is denoted by $|X|$. The image of a matrix $Z\in\mathbb{R}^{n\times m}$ is denoted by $\mathrm{im}\hskip 0.85355ptZ$.
@@ II-A Arbitrary Partition of Nodes
A matrix $S=\mathrm{diag}(s_{1},\ldots,s_{N})\in\mathbb{R}^{N\times N}$ with exactly $1\leq k\leq N$ distinct entries on the diagonal is called a $k$-partition transformation, where each $s_{i}$ is called a $k$-partition transformation term. A 1-partition or 2-partition transformation $S$ is a gauge transformation if $s_{i}\in\{-1,1\}$ for $i=1,\ldots,N$.
Given a signed directed graph, the bipartite framework partitions the nodes according to the signs of the adjacency matrix entries (i.e., the edge weights). This yields at most $2$-partition of nodes. In contrast, we partition the nodes according to a $k$-partition transformation. Since a $k$-partition transformation can be chosen for any $k\in\{1,\ldots,N\}$, the nodes can admit not only 1-partition or 2-partition but also more than 2-partition. In fact, there are infinitely many $k$-partition transformations so that the nodes can be partitioned in the Bell number of ways because $k$-partition transformation terms can be any real number. We now formally discuss the $k$-partition of nodes.
Let $\mathcal{G}(\mathcal{A})$ be a signed directed graph, and let $S$ be a $k$-partition transformation. Then, there exist $k$ positive integers $i_{1},\ldots,i_{k}$ such that $s_{i_{1}},\ldots,s_{i_{k}}$ are $k$ distinct $k$-partition transformation terms. For each $p\in\left\{1,\ldots,k\right\}$, let
@@
(i) Clearly, $i_{p}\in\mathcal{N}_{p}$. (ii) Let $\mathcal{N}_{p}\in\mathcal{C}$ and $\mathcal{N}_{r}\in\mathcal{C}$ with $p\neq r$, but assume for contradiction that $\mathcal{N}_{p}\cap\mathcal{N}_{r}\neq\emptyset$. Then, let $j\in\mathcal{N}_{p}\cap\mathcal{N}_{r}$; hence, $j\in\mathcal{N}_{p}$ and $j\in\mathcal{N}_{r}$. By definition, $s_{j}=s_{i_{p}}$ and $s_{j}=s_{i_{r}}$. Thus, $s_{i_{p}}=s_{i_{r}}$, which contradicts $s_{i_{p}}$ and $s_{i_{r}}$ being distinct. (iii) Since every $\mathcal{N}_{p}\in\mathcal{C}$ is a subset of $\mathcal{N}$, the inclusion $\bigcup_{\mathcal{N}_{p}\in\mathcal{C}}\mathcal{N}_{p}\subseteq\mathcal{N}$ holds. Let $j\in\mathcal{N}$. Then, there exists a $p\in\left\{1,\ldots,k\right\}$ such that $s_{j}=s_{i_{p}}$ due to the fact that $S$ has exactly $k$ distinct diagonal entries. Hence, $j\in\mathcal{N}_{p}$ for some $\mathcal{N}_{p}\in\mathcal{C}$. This proves that the inclusion $\mathcal{N}\subseteq\bigcup_{\mathcal{N}_{p}\in\mathcal{C}}\mathcal{N}_{p}$ holds. ∎
@@
Owing to Definition 2 and Lemma 1, the $k$-partition transformations can achieve arbitrary partition of nodes. For example, let $\mathcal{G}(\mathcal{A})$ be a graph with $5$ nodes. For $k=1,2,3,4,5$, there are infinitely many $k$-partition transformations that can obtain $1,15,25,10,1$ number of $k$-partitions of the nodes, respectively. These numbers correspond to the Stirling numbers of the second kind for respective values of $k$ (e.g., see Section 5.1 in [28]). Their sum is $52$, which is the associated Bell number.
@@ II-B MORP
As in the context of the CORP and BORP, the subsystems of (I-B), considered followers, and the exosystem (2), considered the leader, constitute a MAS of $N+1$ agents. To model the information exchange between $N$ followers, we use a time-invariant signed directed graph ${\mathcal{G}}(\mathcal{A})$ without self-loops as $\mathcal{N}=\left\{1,\ldots,N\right\}$, where node $i\in\mathcal{N}$ corresponds to follower $i$, and for each $j,i\in\mathcal{N}$, we put $(j,i)\in\mathcal{E}$ if, and only if, follower $i$ has access to the information of follower $j$. The leader is included in the information exchange model by augmenting the graph ${\mathcal{G}}(\mathcal{A})$. Specifically, let $\mathcal{G}(\bar{\mathcal{A}})$ be an augmented signed direct graph with $\bar{\mathcal{N}}=\mathcal{N}\cup\{0\}$, $\bar{\mathcal{E}}=\mathcal{E}\cup\mathcal{E}^{\prime}$, where $\mathcal{E}^{\prime}\subseteq\{(0,i)\ |\ i\in\mathcal{N}\}$, and $\bar{\mathcal{A}}\in\mathbb{R}^{(N+1)\times(N+1)}$. Here, node 00 corresponds to the leader and for any $i\in\mathcal{N}$, we put $(0,i)\in\mathcal{E}^{\prime}$ if, and only if, follower $i$ has access to the information of the leader. For any $i\in\mathcal{N}$, the pinning gain $f_{i}>0$ if $(0,i)\in\mathcal{E}^{\prime}$ and $f_{i}=0$ otherwise. The pinning gain matrix is defined by $\mathcal{F}=\mathrm{diag}(f_{1},\ldots,f_{N})$.